### v1

# SounDiT: Geo-Contextual Soundscape-to-Landscape Generation

Abstract
We present a novel and practically significant problem—Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation—which aims to synthesize geographically realistic landscape images from environmental soundscapes. Prior audio-to-image generation methods typically rely on general-purpose datasets and overlook geographic and environmental contexts, resulting in unrealistic images that are misaligned with real-world environmental settings. To address this limitation, we introduce a novel geo-contextual computational framework that explicitly integrates geographic knowledge into multimodal generative modeling. We construct two large-scale geo-contextual multimodal datasets, SoundingSVI and SonicUrban, pairing diverse soundscapes with real-world landscape images. We propose SounDiT, a novel Diffusion Transformer (DiT)-based model that incorporates geo-contextual scene conditioning to synthesize geographically coherent landscape images. Furthermore, we propose a practically-informed geo-contextual evaluation framework, the Place Similarity Score (PSS), across element-, scene-, and human perception-levels to measure consistency between input soundscapes and generated landscape images. Extensive experiments demonstrate that SounDiT outperforms existing baselines in both visual fidelity and geographic settings. Our work not only establishes foundational benchmarks for GeoS2L generation but also highlights the importance of incorporating geographic domain knowledge in advancing multimodal generative models, opening new directions at the intersection of generative AI, geography, urban planning, and environmental sciences.
Figure 1: Geo-contextual soundscape-to-landscape (GeoS2L) generation aims to synthesize realistic landscape images from environmental soundscapes. We introduce large-scale geo-contextual datasets, the SounDiT model, and the Place Similarity Score evaluation framework to support this task.
1 Introduction
“The earth has music for those who listen.” – William Shakespeare
Each place is characterized by a unique set of visual and sounding environments [66, 4, 6, 2], such as natural bird calls in a quiet park starkly contrasting with bustling traffic in downtowns at peak hours. Understanding such geo-contextual linkages regarding how visual and sounding environments co-occur benefits broad applications in geography [74, 14, 39], environmental psychology [53, 26], ecology [60], and urban planning [15]. Integrating these modalities not only advances our knowledge of human-environment interactions [76], but also supports real-world practices [33]. For instance, understanding how the environment shapes human multi-sensory experiences and perceptions [75, 86], informing noise mitigation and place-making practices to benefit human well-being [1, 32], and enhancing inclusive navigation for visually impaired populations [37, 8].
Prior research in the Audio-to-Image (audio2image) stream that translates auditory information into visual representations has offered promising opportunities [70, 61, 80, 79, 5]. However, we observed two noticeable gaps when applying prior studies in geographic contexts and planning practices. First, most audio2image studies rely on generic audio-visual datasets such as object audio [57, 85, 9], human voice [54, 49], weathers [42], and limited scene types [41], and overlook geo-contextual insights such as parks vs. beaches and urban vs. rural settings [81] (See Appendix LABEL:appendix:dataset). Geographers are primarily interested in the environmental context of a soundscape rather than identifying sound sources such as objects or animals. Second, common evaluation metrics employed in prior studies, such as the Fréchet Inception Distance (FID) [30], may fall short in capturing geographic semantic alignment between the input soundscape and generated images. Ideally, the evaluation metrics should consider environmental characteristics and geographic contexts: the place setting depicted in the input soundscape and those represented in the generated images should be similar to ensure geographic contextual coherence and relevance. Furthermore, Diffusion Transformer (DiT) [59] techniques have recently demonstrated impressive results across various tasks such as image [44, 10, 11] and video generations [51, 77, 38, 48]. However, they remain unexplored in audio2image generation and have yet to incorporate geographic knowledge to support applications in geography and urban planning.
Given these challenges, we introduce and formalize a new problem, namely, Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation (Figure 1), to synthesize realistic landscape images from environmental soundscapes, optionally conditioned on scene context. The GeoS2L problem extends traditional audio2image synthesis by incorporating geo-contextual insights with significant practical implications. Here, landscape refers to geographical environments shaped by natural and built features [52]. A soundscape denotes the acoustic environment as perceived by humans at places, comprising sounds from both natural and anthropogenic sources (e.g., bird calls and vehicle noises) [68]. A scene indicates place types such as beach, forest, street, to guide image generation. Our contributions span three key areas: First, we introduce two large-scale, multi-modal geo-contextual datasets: SoundingSVI and SonicUrban, composed of over 330K and 250K soundscape-landscape image pairs. These datasets cover diverse geographic locations and scenes, supporting geo-contextual realistic synthesis. Second, we develop an innovative generative model, SounDiT, which integrates explicit geographic contexts into DiT to effectively translate environmental soundscapes into corresponding geographically coherent landscape images. Third, we design a novel geo-contextual evaluation framework, namely, Place Similarity Scores (PSS), which includes three metrics: element-level, scene-level, and perception-level similarity to assess geographic contexts and environmental consistency of generated landscape images. Our proposed SounDiT achieves state-of-the-art performance on SoundingSVI and SonicUrban. Beyond technical contributions, this work supports applications in geography, environment, and urban planning with significant practical impacts.
2 Related Work
Audio-to-Image Synthesis aims to translate acoustic features into coherent and meaningful images and has gained increasing attention with the advancement of various multi-modal alignment foundation models [79, 70] and generative techniques [24, 65]. Current audio2image models are broadly divided into GAN-based (e.g., sound2scene [70]) and Diffusion-based (e.g., AudioToken [80], and GluenGen [61]). These models mainly rely on generic datasets which lack geo-contextual information, often limited to object audios [57], human voice [54], weather [42], and broad scene types [41]. As a result, they often produce stylized or unrealistic outputs and fail to capture the geographic coherence between soundscapes and landscapes (See Appendix LABEL:appendix:dataset).
Landscapes and Soundscapes represent two dimensions for characterizing places [46, 19]. Landscapes offer visual representations of the physical layout and built environments, and are commonly captured via street view imagery (SVI) and remote sensing (RS) images [41, 42]. In parallel, soundscapes refer to the acoustic environment perceived at specific places. Existing datasets, such as BirdCLEF [34], BirdSet [64], SoundingEarth [28]), AudioSet [20], UrbanSound8K [67], SONYC [7], EnigenScape [25], and EMO-Soundscape [16], have been utilized to support tasks including monitoring biodiversity [63], noise pollution [32], and human perception analysis [18]. However, prior audio2image studies have overlooked such geographic and environmental insights, limiting their relevance for real-world applications. To address this gap, our work integrates geo-contextual information for multimodal generative modeling by constructing datasets and evaluation frameworks.
3 Method
3.1 Preliminary
Diffusion Model. We adopt a latent-based diffusion model, consisting of a forward diffusion and backward denoising process conducted in the latent space $\mathbf{z}$. The forward diffusion $\mathbf{q}$ starts from the initial latent $\mathbf{z}_{0}$, gradually adding Gaussian noise over timestep $T$ based on a fixed variance $\alpha$ and defined as:
$q(\mathbf{z}_{t}\mid\mathbf{z}_{t-1})=\mathcal{N}\bigl{(}\mathbf{z}_{t};\sqrt{\alpha_{t}}\,\mathbf{z}_{t-1},(1-\alpha_{t})\,\mathbf{I}\bigr{)},\quad t=1,\dots,T.$
The backward denoising process $p_{\theta}$ is trained to predict the less noisy latent:
$p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t})=\mathcal{N}\bigl{(}\mathbf{z}_{t-1};\,\mu_{\theta}(\mathbf{z}_{t},t,v),\,\Sigma_{\theta}(\mathbf{z}_{t},t,v)\bigr{)},$
where $v$ is the conditional signal and $\mu_{\theta}$, $\Sigma_{\theta}$ are predicted through a denoising network $\epsilon_{\theta}(\mathbf{z}_{t},t,v)$.
Classifier-Free Guidance. To control the influence of the conditioning signal $v$ during backward denoising, we employ classifier-free guidance (CFG) to predict noise $\hat{\boldsymbol{\epsilon}}$ based on both conditional and unconditional signals with specific guidance scale $s$.
$\hat{\boldsymbol{\epsilon}}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)+s\bigl{[}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,v)-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)\bigr{]}.$
3.2 Geo-Contextual Soundscape-to-Landscape Generation (GeoS2L) Problem Definition
In this work, we formulate our task as the Geo-Contextual Soundscape-to-Landscape (GeoS2L) problem, a specialized variant of the audio2image problem, which aims to generate landscape images from environmental soundscapes guided by an optional scene prompt. We define a landscape $l$ as the geographical and ecological environment shaped by natural and built features, and a soundscape $s$ as the human-perceived acoustic environment at a given place, comprising both natural and anthropogenic sounds such as bird calls and traffic noise. To better support real-world practices, we introduce scene prompt $c$ as an optional semantic label (e.g., park, beach, street) to provide additional geographic semantic context.
Given a dataset $D=\{\,(s_{i},c_{i},l_{i})\mid i=1,\dots,N\},$ our goal is to learn a generator $G\colon(s_{i},c_{i})\mapsto\hat{l}_{i}$ such that the synthesized image $\hat{l}_{i}$ is both visually realistic and geo-contextually consistent to the ground truth $l_{i}$, quantified by relevance function $R(s,c,l)$. This can be formulated as:
$\mathcal{L}\;=\;\mathbb{E}_{(s_{i},c_{i},l_{i})\sim D}\Bigl{[}\,R(s_{i},c_{i},l_{i})\;-\;R\bigl{(}s_{i},c_{i},\hat{l}_{i}\bigr{)}\Bigr{]}.$
Minimizing $\mathcal{L}$ enables the model to capture cross-modal correlations among sounding, semantic and visual features, thereby producing landscape images that reflect both the acoustic characteristics of soundscape and intended scene.
Figure 2: Framework of SounDiT. To effectively capture geographic contexts, SounDiT integrates soundscape, scene, and landscape embeddings from pre-trained encoders as inputs. The landscape embedding is fused with the timestep embedding and fed into a self-attention layer to extract visual features. The scene embedding then guides conditioning via a cross-attention layer. The TA-PMoE module further combines these features with soundscape features. During inference, a geo-contextual RAG module provides additional geographic contexts to enhance alignment with real-world environmental settings.
3.3 SounDiT
To support the GeoS2L task, we propose SounDiT, a diffusion-based generative model that synthesizes geographically coherent landscape images conditioned on environmental soundscapes and optional scene prompts (Figure 2). SounDiT advances geo-contextual landscape generation through two key innovations: (1) the SounDiT Block, which integrates sound, scene, and visual cues into the denoising process; and (2) a geo-contextual Retrieval-Augmented Generation (RAG) to guide image generation with additional geographic contexts at inference time.
Multimodal Encoders. To incorporate acoustic, semantic and visual features into SounDiT, we employ three encoder modules: (1) a Soundscape Encoder $\mathcal{E}_{\mathrm{Sound}}$, (2) a Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ and (3) a Landscape Encoder–Decoder pair $\bigl{(}\mathcal{E}_{L},\mathcal{E}^{\prime}_{L}\bigr{)}$. The Soundscape Encoder $\mathcal{E}_{Sound}$ embeds soundscape $s$ into an audio embedding $e_{\mathrm{sound}}$. The Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ encodes the optional scene prompt $c$ into a text embedding $e_{\mathrm{scene}}$. The Landscape Encoder $\mathcal{E}_{L}$ compresses a landscape image $l_{i}\in\mathbb{R}^{H\times W\times 3}$ into a latent embedding $e_{\mathrm{landscape}}$, and the Landscape Decoder $\mathcal{E^{\prime}}_{L}$ reconstructs the latent embedding back into a landscape image $\hat{l_{i}}\in\mathbb{R}^{H\times W\times 3}$.
SounDiT Block. Despite the demonstrated effectiveness of DiT in image generation, they have not been integrated into audio2image tasks. To effectively model geographic contexts from both soundscape and scene inputs, we propose the SounDiT block, a geo-contextual DiT to generate landscape images that align with intended geographic settings. The SounDiT operates in three stages: First, the timestep embedding $e_{\mathrm{t}}$ is fused with landscape latent embedding $e_{\mathrm{landscape}}$, and fed into a multi-head self-attention layer to capture visual features. Second, a multi-head cross-attention conditioned on the scene embedding $e_{\mathrm{scene}}$ is applied to enrich these features with scene semantics, producing scene-landscape representation. Third, to further integrate multi-level soundscape features, we introduce Time-Aware Prototype Mixture-of-Experts (TA-PMoE) module, composed of $M$ parallel cross-attention layers as experts. Each cross-attention layer fuses the scene-landscape representation with the soundscape embedding $e_{\mathrm{sound}}$, resulting in $M$ soundscape-scene-landscape representations $\mathbf{P}=[\,\mathbf{p}_{1},\dots,\mathbf{p}_{M}\,]$. A router $\mathbf{w}$ then calculates a weighted combination of these representations $\mathbf{P}$ based on temperature-scaled Softmax:
[…]
Here, $\tau>0$ controls the sharpness of the distribution: smaller values of $\tau$ concentrate weight on the most relevant soundscape-scene-landscape representation, while larger values lead to a smoother allocation across representations. Finally, the combined soundscape-scene-landscape representation is passed through a pointwise feedforward network to produce the denoised output. By hierarchically fusing visual, scene, and soundscape representations, SoundDiT effectively generates geo-contextually coherent landscape images.
Geo-contextual RAG Inference. In practice, additional geo-contextual information, such as location, scene and spatial layout [47], may be collected alongside soundscapes and landscapes to guide image generation. Thus, we propose a flexible geo-contextual RAG pipeline to further enhance the geographic coherence of the generated landscape images. Our method consists of two stages: (1) Geo-contextual Knowledge Database (GeoKD), where we store soundscape embeddings along with associated scene categories and layout representations with geographic locations; and (2) Geo-contextual Initialization, where the most relevant scene and layout are retrieved based on soundscape similarity and geographic proximity. These retrieved cues are then used to initialize the diffusion generation to enhance geographic coherence.
4 Datasets: SoundingSVI and SonicUrban
Prior audio2image studies have primarily relied on general datasets such as VEGAS [85], VGGSound [9], and ACAV100M [40]. Despite their proven effectiveness, these datasets lack geographic contexts. For instance, geographers may be more interested in generating a wetland implied by a frog call, rather than an image of the frog itself. A car horn might suggest a congested urban street, rather than a static vehicle. Some datasets, such as Landscape [41] and IntoTheWild [42], have included paired soundscape and image data, but offer limited spatial coverage and scene diversity. To address this gap, we introduce two new multimodal datasets: SoundingSVI and SonicUrban, designed to support GeoS2L generation with large-scale, geographically diverse, and context-rich data. Figure 3 provides several sample images of these two datasets.
[…]
Dataset Modalities # Size Geo-Contextual Scene Diversity Spatial Coverage Viewpoint
ACAV100M [40] Video 100M -- -- -- Ground-level
VEGAS [85] Video 28K -- -- -- Ground-level
[…]
IntoTheWild [42] Video 142 $\checkmark$ Low -- Ground-level
Landscape [41] Video 1747 $\checkmark$ Medium -- Ground-level
SoundingEarth [28] Audio & Image 50K $\checkmark$ High High Overhead
SoundingSVI (Ours) Audio & Image 330K $\checkmark$ High High Ground-level
SonicUrban (Ours) Video 250K $\checkmark$ High High Ground-level
Figure 3: Representative samples from the SoundingSVI and SonicUrban datasets, showing diverse geographic contexts.
SoundingSVI. The SoundingSVI dataset was constructed from geotagged soundscape recordings from the Aporee platform [56]. For each raw soundscape recording, we retrieved nearby Google Street View Image (SVI) using its latitude and longitude to capture the surrounding visual landscapes. This follows a similar approach to the Sounding Earth dataset [28], which paired geotagged audio with remote sensing imagery. The raw recordings were then segmented into 10-second clips, with a human voice detection model [73] used to filter clips dominated by human voice. Notably, multiple candidate SVIs captured from different angles and timestamps might be retrieved for clips from the same recording. A sound source localization model [58] was further applied to match each clip with the most relevant landscape image. In total, SoundingSVI dataset comprises 334,495 soundscape-landscape pairs in 90 countries. To support optional scene prompts during training, a scene type for each soundscape-landscape pair was predicted by a scene classification model [27].
SonicUrban. The SonicUrban dataset leverages unedited videos as a source of temporally and spatially consistent soundscape-landscape pairs. Following a prior practice [86], we searched and manually checked YouTube videos that contain real-world soundscapes and landscapes, using a combination of city names and keywords, like “City walk; New York”. Each video was processed into 10-second audio clips, and 10 evenly distributed frames were extracted for each clip. The same human voice detection model [73] and sound source localization model [58] used in SoundingSVI were applied to filter human-voice dominated clips and identify the most representative frame for each audio clip, respectively. In total, the constructed SonicUrban consists of 250,310 soundscape-landscape pairs across 131 cities and 97 countries. A scene classification model [27] was further employed to generate a scene prompt for each pair as a geographic scene condition.
5 Geo-Contextual Evaluation Framework
Figure 4: Geo-contextual evaluation framework. The ground truth and output landscape images are assessed at the element-, scene-, and human-perception levels following geography and urban planning practices.
Prior audio2image studies commonly rely on evaluation metrics such as Fréchet Inception Distance (FID) [30], Audio-Image Similarity (AIS) [30], and Image-Image Similarity (IIS) [5], to assess visual or audio fidelity, overlooking the underlying geographic settings and target scenes. Since soundscapes and landscapes co-occur in the same space, they are expected to share similar environmental characteristics and place settings. Following this hypothesis, we introduce a novel practically-informed geo-contextual evaluation framework — Place Similarity Score (PSS) — that measures the environmental settings of soundscape and landscape across element, scene, and human perception levels (Figure 4). These three levels of measurements of environmental characteristics have been widely employed in geographic and urban planning practices [50, 36].
Table 2: Place Similarity Score (PSS)
Score Definition
$\text{PSS}_{\text{element}}$ $\frac{1}{n}\sum_{i=1}^{n}\frac{E_{i}\cdot\hat{E}_{i}}{\|E_{i}\|\|\hat{E}_{i}\|}$
$\text{PSS}_{\text{scene}}$ $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{P_{i}^{\text{k}}\cap T_{i}^{\text{k}}\neq\varnothing\right\}$
$\text{PSS}_{\text{perception}}$ $\frac{1}{n}\sum_{i=1}^{n}\left\|R(l_{i})-R(\hat{l}_{i})\right\|_{1}$
Element-level Evaluation quantifies the similarity of geographic elements (e.g., trees [43], sky [23], water bodies [69], traffic signs [3], and buildings [13]) between the geographical contexts of the input soundscape and generated landscapes [35]. We employ the DeepLabV3 semantic segmentation model [12], pre-trained on the ADE20K dataset [84], to identify 150 predefined geographic elements $E_{i}$ in each image $i$. For each image $i$, we calculate a normalized element distribution vector $E_{i}=\{e_{i}^{1},e_{i}^{2},\dots,e_{i}^{k}\}$, where each $e_{i}^{k}$ represents the ratio of pixels identified as the $k$-th element. [17]. We calculate element ratio vectors $E_{i}$ and $\hat{E_{i}}$ from the ground truth landscape image $l_{i}$ and their corresponding image $\hat{l_{i}}$. We then measure the cosine similarity between each pair of real and generated images and report the average across all samples $n$ as the $PSS_{element}$.
Scene-level Evaluation is designed to evaluate the consistency of the entire environmental scene (e.g., “forest,” “beach,” or “residential area”) between the generated landscape images and their corresponding ground truth images [35]. To infer scene categories, we employ a ResNet50 model [27], pre-trained on the Places365 dataset [83], to identify 365 predefined scene categories. For each soundscape $s$, we predict the top-$k$ scene labels and evaluate whether the generated image $\hat{l_{i}}$ aligns with the same scene category as its ground truth image $l_{i}$. The $PSS_{scene}$ is then calculated by its top-$k$ accuracy, where $k=1$ or $5$. $n$ is the total number of images, $P_{i}^{\text{k}}$ denotes the top-$k$ predicted category set for the $i$-th generated image, $T_{i}^{\text{k}}$ denotes the top-$k$ ground truth category set for the $i$-th real image.
Human Perception-level Evaluation is designed to evaluate whether the generated and real landscape images evoke similar human subjective perceptions and feelings, such as whether a place makes people feel “safe” or “lively” [82]. We utilize a DenseNet121 model [31], pre-trained on the MIT Place Pulse dataset [55], to measure six dimensions of human perceptions of environment, including “Safe,” “Beautiful,” “Depressing,” “Lively,” “Wealthy,” and “Boring”. For each image $i$, we compute the perception scores in six dimensions for the generated image $\mathrm{R}(\hat{l}_{i})$ and its corresponding ground truth $\mathrm{R}(l_{i})$. $PSS_{perception}$ is then defined as the average L1 distance across six dimensions, where $n$ is the number of the test set, $R$ is the Place Perception scores of each dimension.
By integrating metrics across element-, scene-, and human perception levels (Table 2), we provide a holistic geo-contextual evaluation framework. Informed by geographic domain knowledge and urban planning practices, we extend beyond measuring visual quality to assessing whether generated landscape images are contextually aligned with the environmental characteristics in input soundscapes.
6 Experiments
Figure 5: Visual comparison of landscape images generated by baseline models (CoDi, Sound2Scene, AudioToken Pre-trained/SD1/SD2, GlueGen) and our proposed SounDiT across SoundingSVI, SonicUrban, and two benchmark datasets. Ground truth landscape images are provided in the last column. Our proposed SounDiT captures significantly more geographic contexts than baselines.
6.1 Implementation Details
Model Architecture and Training. We adopt the pre-trained Variational Autoencoder (VAE) modules from Latent Stable Diffusion [65] as our landscape encoder–decoder pair, which were originally trained on the COCO dataset [45]. For soundscape conditioning, we employ the ImageBind-Huge audio encoder [22], pre-trained on two million AudioSet clips [21] covering a wide range of environmental soundscapes. We integrate a scene encoder based on the T5 architecture [62], which transforms high-level semantic scene prompts into latent embeddings to provide geographic semantics. Our SounDiT Block is initialized with the weights of PixArt-$\alpha$, a pre-trained text-to-image DiT model [10], and four experts are employed in TA-MoE Module. All components are trained using a learning rate of $1\times 10^{-4}$. Experiments are conducted on NVIDIA A100 and A6000 GPUs.
Benchmark Datasets. We curated two additional GeoS2L benchmark datasets to evaluate model generalizability. The first is a geo-contextual subset of VGGSound [9], comprising 16 classes with environmental semantics. It contains 6,713 soundscape–landscape pairs. The second combines the Landscape [41] and IntoTheWild [42] datasets, resulting in 1,889 soundscape-landscape pairs.
Baseline Models. We selected four state-of-the-art audio2image models, including CoDi [72], Sound2Scene [70], AudioToken [80], and GlueGen [61] as baseline models. For CoDi, we utilized its official pre-trained model. For AudioToken, we employed the pre-trained model without additional training, and trained it on our datasets using Stable Diffusion 1 and 2 as its backbone, respectively. The other two baselines are training from scratch based on their released code and training parameters.
Benchmark Evaluation Metrics. To evaluate the model performance, we adopted FID [30], AIS [30], and IIS [5], effective in previous audio-conditioned image generation studies [5, 71]. FID compares Gaussian distributions of features extracted from real and generated images, using a pre-trained Inception model. AIS measures audio-image semantic similarity using Wav2CLIP [78]. IIS evaluates image-image semantic similarity using CLIP-based embedding [29].
6.2 Main Results
[…]
General Metrics Place Similarity Score
[…]
Sound2Scene 78.137 0.586 0.623 0.467 0.374 0.735
[…]
AudioToken (SD1) 105.472 0.509 0.572 0.527 0.553 0.797
AudioToken (SD2) 149.473 0.504 0.559 0.249 0.311 0.845
GlueGen 74.250 0.500 0.566 0.400 0.461 0.776
SounDiT (Ours) 28.173 0.542 0.663 0.652 0.578 0.728
[…]
Sound2Scene 49.681 0.503 0.498 0.332 0.324 0.829
[…]
AudioToken (SD1) 95.210 0.505 0.503 0.419 0.484 0.979
AudioToken (SD2) 165.014 0.502 0.504 0.298 0.455 0.980
GlueGen 64.434 0.498 0.499 0.405 0.481 0.805
SounDiT (Ours) 33.246 0.517 0.592 0.465 0.495 0.788
[…]
Sound2Scene 140.244 0.617 0.741 0.479 0.640 0.769
[…]
AudioToken (SD1) 138.785 0.585 0.695 0.364 0.534 0.846
AudioToken (SD2) 118.536 0.637 0.740 0.512 0.724 0.855
GlueGen 195.625 0.584 0.645 0.362 0.582 0.815
SounDiT (Ours) 155.471 0.628 0.744 0.521 0.741 0.757
[…]
Sound2Scene 122.460 0.600 0.596 0.287 0.378 0.770
[…]
AudioToken (SD1) 163.540 0.532 0.555 0.207 0.311 0.789
AudioToken (SD2) 157.706 0.583 0.587 0.251 0.404 0.803
GlueGen 156.777 0.536 0.547 0.233 0.249 0.839
SounDiT (Ours) 113.111 0.575 0.588 0.336 0.392 0.767
Figure 5 presents landscape images generated by our proposed SounDiT and baseline models across SoundingSVI, Sonic Urban, and two benchmark datasets. The results generated by SounDiT show greater geo-contextual similarity to ground truth images, highlighting potential applications of SounDiT in geography and urban planning.
To evaluate model performance, we adopt general metrics (FID, AIS, IIS) in addition to our proposed geo-contextual PSS framework. As reported in Table 3, SounDiT consistently outperforms baselines on both SoundingSVI and SonicUrban datasets, with notable FID improvements (74$\rightarrow$28 and 49$\rightarrow$33), and achieves superior or competitive results across all metrics. SounDiT consistently ranks highest under PSS, indicating its ability to ensure geographic coherence in generated images. These results validate the effectiveness of PSS in measuring geo-contextual coherence beyond general metrics. The performance drop of SounDiT on the smaller IntoTheWild&Landscape dataset further emphasizes the crucial needs for large-scale training datasets such as SoundingSVI and SonicUrban.
User Study. In addition to these quantitative metrics, we further conducted a user study to assess whether the generated landscape images align with human perceptions of geographic contexts. Specifically, participants were asked to complete two matching tasks: (1) selecting the generated image that best aligned with the given soundscape; and (2) identifying which generated landscape image most closely matched a ground-truth landscape. The average matching accuracy is 86.13%, indicating a strong perceptual alignment between soundscapes and their corresponding generated images. This demonstrates the effectiveness of SounDiT for GeoS2L.
6.3 Ablations
Table 4: Ablation results. We assess the contribution of key components to the overall performance of SounDiT.
[…]
TA-MoE 28.173 0.652
MHCA 34.108 0.602
(a) SounDiT Block Design
[…]
w/ RAG 0.652 0.578
w/o RAG 0.479 0.514
(b) Geo-contextual RAG
We conducted the ablation study on SoundingSVI. First, we replace the TA-PMoE module with a standard multi-head cross-attention layer, forming SounDiT-MHCA. Second, we disable the RAG module in SounDiT-TA-MoE, yielding SounDiT-TA-MoE w/o RAG. As reported in Table 4(a), SounDiT-TA-MoE achieves a lower FID and a higher $PSS_{element}$, indicating that the expert routing (TA-MoE) is essential for modeling multi-level soundscape and visual features, thereby preserving visual quality and geographic coherence in the generated images. Table 4(b) further shows that incorporating geo-contextual RAG inference improves both element- and scene-level PSS of generated landscape images, highlighting the benefits of integrating geographic contexts during inference.
7 Concluding Remarks
In this work, we present Geo-Contextual Soundscape-to-Landscape (GeoS2L), a practically-informed novel generation problem that synthesizes realistic landscape images from environmental soundscapes. We explicitly integrate geographic domain knowledge into multimodal generative modeling. To support GeoS2L, we construct SoundingSVI and SonicUrban, two large-scale, multi-modal, geo-contextual datasets, capturing diverse sounding and visual environments. We propose SounDiT, a novel diffusion-based framework that captures geographic semantics embedded in sounding environments. We further design a novel geo-contextual evaluation framework, the Place Similarity Score (PSS), which assesses geographic and environmental consistency. Extensive experiments demonstrate that SounDiT consistently outperforms existing baselines and can generate realistic landscape images. The proposed geo-contextual datasets, model, and evaluation framework effectively integrate geographic knowledge into generative modeling, establishing a solid benchmark for the GeoS2L task.
Limitations. While our work introduces a novel problem, GeoS2L, several limitations remain. First, the SoundingSVI dataset relies on geotagged audio from Aporee, which is a crowd-sourced platform that may contain spatial inaccuracies. Future efforts may focus on refining spatial alignment and improving dataset quality through more robust data processing procedures. Second, the current PSS evaluation metrics only consider three practically-informed levels. Future studies could integrate more geographic domain-specific knowledge to support geo-contextually coherent image generation.

### v2

@@
-SounDiT: Geo-Contextual Soundscape-to-Landscape Generation
+arXiv:2505.12734v2[cs.SD] 27 Feb 2026
 Abstract
-We present a novel and practically significant problem—Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation—which aims to synthesize geographically realistic landscape images from environmental soundscapes.
-Prior audio-to-image generation methods typically rely on general-purpose datasets and overlook geographic and environmental contexts, resulting in unrealistic images that are misaligned with real-world environmental settings.
-To address this limitation, we introduce a novel geo-contextual computational framework that explicitly integrates geographic knowledge into multimodal generative modeling.
-We construct two large-scale geo-contextual multimodal datasets, SoundingSVI and SonicUrban, pairing diverse soundscapes with real-world landscape images.
-We propose SounDiT, a novel Diffusion Transformer (DiT)-based model that incorporates geo-contextual scene conditioning to synthesize geographically coherent landscape images.
-Furthermore, we propose a practically-informed geo-contextual evaluation framework, the Place Similarity Score (PSS), across element-, scene-, and human perception-levels to measure consistency between input soundscapes and generated landscape images.
-Extensive experiments demonstrate that SounDiT outperforms existing baselines in both visual fidelity and geographic settings.
-Our work not only establishes foundational benchmarks for GeoS2L generation but also highlights the importance of incorporating geographic domain knowledge in advancing multimodal generative models, opening new directions at the intersection of generative AI, geography, urban planning, and environmental sciences.
+Recent audio-to-image models have shown impressive performance in generating images of specific objects conditioned on their corresponding sounds.
+However, these models fail to reconstruct real-world landscapes conditioned on environmental soundscapes.
+To address this gap, we present Geo-contextual Soundscape-to-Landscape (GeoS2L) generation, a novel and practically significant task that aims to synthesize geographically realistic landscape images from environmental soundscapes.
+To support this task, we construct two large-scale geo-contextual multi-modal datasets, SoundingSVI and SonicUrban, which pair diverse environmental soundscapes with real-world landscape images.
+We propose SounDiT, a diffusion transformer (DiT)-based model that incorporates environmental soundscapes and geo-contextual scene conditioning to synthesize geographically coherent landscape images.
+Furthermore, we propose the Place Similarity Score (PSS), a practically-informed geo-contextual evaluation framework to measure consistency between input soundscapes and generated landscape images.
+Extensive experiments demonstrate that SounDiT outperforms existing baselines in the GeoS2L, while the PSS effectively captures multi-level generation consistency across element, scene,and human perception.
+Project page: https://gisense.github.io/SounDiT-Page/
 Figure 1: Geo-contextual soundscape-to-landscape (GeoS2L) generation aims to synthesize realistic landscape images from environmental soundscapes.
@@
 1 Introduction
-“The earth has music for those who listen.” – William Shakespeare
-Each place is characterized by a unique set of visual and sounding environments [66, 4, 6, 2], such as natural bird calls in a quiet park starkly contrasting with bustling traffic in downtowns at peak hours.
-Understanding such geo-contextual linkages regarding how visual and sounding environments co-occur benefits broad applications in geography [74, 14, 39], environmental psychology [53, 26], ecology [60], and urban planning [15].
-Integrating these modalities not only advances our knowledge of human-environment interactions [76], but also supports real-world practices [33].
-For instance, understanding how the environment shapes human multi-sensory experiences and perceptions [75, 86], informing noise mitigation and place-making practices to benefit human well-being [1, 32], and enhancing inclusive navigation for visually impaired populations [37, 8].
-Prior research in the Audio-to-Image (audio2image) stream that translates auditory information into visual representations has offered promising opportunities [70, 61, 80, 79, 5].
-However, we observed two noticeable gaps when applying prior studies in geographic contexts and planning practices.
-First, most audio2image studies rely on generic audio-visual datasets such as object audio [57, 85, 9], human voice [54, 49], weathers [42], and limited scene types [41], and overlook geo-contextual insights such as parks vs.
+“The earth has music for those who listen.”
+— William Shakespeare
+Environmental soundscapes encode rich geographic information, often in greater detail than the specific object acting as the sound source [67, 4, 6, 1].
+A burst of birdsong, for example, implies a leafy trail or an urban green space, while the sound of traffic may evoke a busy streetscape, beyond a single car.
+Recognizing this strong geo-contextual connection between auditory soundscape and visual landscape, researchers in geography [75, 16, 43], urban planning [17, 39, 31], and environmental psychology [56, 28], have long studied the impacts of soundscapes.
+For example, how analyzing specific sound types can inform and improve urban design[36], how to mitigate noise through environmental design [7, 8], or which soundscapes are associated with human restoration and well-being [84, 76].
+Previous practices in these fields have often been limited to descriptive statistics of the soundscape environment (e.g., decibel levels, sound types) [37, 15], which are not always intuitive and fail to capture the co-located visual characteristics of the scene.
+The recent emergence of deep generative models for Audio-to-Image (A2I) synthesis [72, 63, 81, 80, 5], which translates auditory information directly into visual representations, offers a promising opportunity to model the geo-contextual soundscape-landscape relationships, unlocking significant practical applications.
+However, we observe three noticeable gaps when applying prior A2I in geographic contexts and planning practices for associating soundscapes and landscapes.
+First, most A2I studies rely on generic audio-visual datasets such as object audio [60, 87, 10], human voice [57, 52], weathers [46], and limited scene types [45].
+Consequently, these models link an audio signal to its source (e.g., generating an image of a bird for a bird call) rather than its environment, overlooking geo-contextual insights vital for real-world applications, such as parks vs.
 beaches and urban vs.
-rural settings [81] (See Appendix LABEL:appendix:dataset).
-Geographers are primarily interested in the environmental context of a soundscape rather than identifying sound sources such as objects or animals.
-Second, common evaluation metrics employed in prior studies, such as the Fréchet Inception Distance (FID) [30], may fall short in capturing geographic semantic alignment between the input soundscape and generated images.
+rural settings [82].
+For geographers and urban planners, the primary interest is not the object making the sound (that bird, that car), but the environmental setting that informs environmental design and noise mitigation strategies.
+Despite explorations of diffusion-based models in geographic audio-to-image settings [89, 88], these methods still take soundscapes as the sole input and output landscape images, without geographic context.
+Second, Diffusion Transformer (DiT) [62] techniques have recently demonstrated impressive results across various tasks such as image [48, 12, 11] and video generations [54, 77, 42, 51].
+However, they remain unexplored in A2I generation and have yet to incorporate geographic knowledge to support applications in geography and urban planning.
+Furthermore, common evaluation metrics employed in prior studies, such as the Fréchet Inception Distance (FID) [33], may fall short in capturing geographic semantic alignment between the input soundscape and generated images.
 Ideally, the evaluation metrics should consider environmental characteristics and geographic contexts: the place setting depicted in the input soundscape and those represented in the generated images should be similar to ensure geographic contextual coherence and relevance.
-Furthermore, Diffusion Transformer (DiT) [59] techniques have recently demonstrated impressive results across various tasks such as image [44, 10, 11] and video generations [51, 77, 38, 48].
-However, they remain unexplored in audio2image generation and have yet to incorporate geographic knowledge to support applications in geography and urban planning.
-Given these challenges, we introduce and formalize a new problem, namely, Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation (Figure 1), to synthesize realistic landscape images from environmental soundscapes, optionally conditioned on scene context.
-The GeoS2L problem extends traditional audio2image synthesis by incorporating geo-contextual insights with significant practical implications.
+Given these challenges, we introduce and formalize a new, practically-motivated task: Geo-Contextual Soundscape-to-Landscape (GeoS2L) generation (Figure SounDiT: Geo-Contextual Soundscape-to-Landscape Generation), to synthesize realistic landscape images from environmental soundscapes, optionally conditioned on scene context.
+The GeoS2L problem extends traditional A2I synthesis by incorporating geo-contextual insights with significant practical implications.
 Here, landscape refers to geographical environments shaped by natural and built features [55].
@@
 A scene indicates place types such as beach, forest, street, to guide image generation.
-Our contributions span three key areas: First, we introduce two large-scale, multi-modal geo-contextual datasets: SoundingSVI and SonicUrban, composed of over 330K and 250K soundscape-landscape image pairs.
+Our contributions span three key areas: First, we introduce two large-scale, multi-modal geo-contextual datasets: SoundingSVI and SonicUrban, composed of over 169K and 237K soundscape-landscape image pairs.
 These datasets cover diverse geographic locations and scenes, supporting geo-contextual realistic synthesis.
-Second, we develop an innovative generative model, SounDiT, which integrates explicit geographic contexts into DiT to effectively translate environmental soundscapes into corresponding geographically coherent landscape images.
-Third, we design a novel geo-contextual evaluation framework, namely, Place Similarity Scores (PSS), which includes three metrics: element-level, scene-level, and perception-level similarity to assess geographic contexts and environmental consistency of generated landscape images.
-Our proposed SounDiT achieves state-of-the-art performance on SoundingSVI and SonicUrban.
-Beyond technical contributions, this work supports applications in geography, environment, and urban planning with significant practical impacts.
+Second, we develop SounDiT, an innovative generative model that support GeoS2L problem.
+It introduces three novel, efficient components to integrate geo-contextual information: a MoE Soundscape Conditioning module, a Scene Low-Rank Content Mixer (SLRCM), and a Scene AdaLn (S-AdaLn).
+These modules effectively inject geographic contexts to environmental soundscapes, enabling the generation of geographically coherent landscape images.
+Third, we design a novel practically-motivated geo-contextual evaluation framework: Place Similarity Scores (PSS).
+PSS includes three metrics: element-level, scene-level, and perception-level similarity to assess geographic contexts and environmental consistency of generated landscape images.
+Extensive experiments demonstrate that our model achieves state-of-the-art performance and delivers strong utility on downstream applications like geography, environment, and urban planning with significant practical impacts.
+Moreover, our datasets and evaluation metric provide a robust benchmark for the GeoS2L task.
 2 Related Work
 Audio-to-Image Synthesis aims to translate acoustic features into coherent and meaningful images and has gained increasing attention with the advancement of various multi-modal alignment foundation models [80, 72] and generative techniques [25, 66].
-Current audio2image models are broadly divided into GAN-based (e.g., sound2scene [70]) and Diffusion-based (e.g., AudioToken [80], and GluenGen [61]).
+Current A2I models are broadly divided into GAN-based (e.g., sound2scene [72]) and Diffusion-based (e.g., AudioToken [81], and GluenGen [63]).
 These models mainly rely on generic datasets which lack geo-contextual information, often limited to object audios [60], human voice [57], weather [46], and broad scene types [45].
-As a result, they often produce stylized or unrealistic outputs and fail to capture the geographic coherence between soundscapes and landscapes (See Appendix LABEL:appendix:dataset).
+As a result, they often produce stylized or unrealistic outputs and fail to capture the geographic coherence between soundscapes and landscapes.
 Landscapes and Soundscapes represent two dimensions for characterizing places [50, 20].
@@
 In parallel, soundscapes refer to the acoustic environment perceived at specific places.
-Existing datasets, such as BirdCLEF [34], BirdSet [64], SoundingEarth [28]), AudioSet [20], UrbanSound8K [67], SONYC [7], EnigenScape [25], and EMO-Soundscape [16], have been utilized to support tasks including monitoring biodiversity [63], noise pollution [32], and human perception analysis [18].
-However, prior audio2image studies have overlooked such geographic and environmental insights, limiting their relevance for real-world applications.
+Existing datasets, such as BirdCLEF [38], BirdSet [65], SoundingEarth [30]), AudioSet [22], UrbanSound8K [68], SONYC [9], EnigenScape [27], and EMO-Soundscape [18], have been utilized to support real-world practices including monitoring biodiversity [64], noise mitigation [35], and human perception analysis [19].
+However, prior A2I studies have overlooked such geographic and environmental insights, limiting their real-world applications.
 To address this gap, our work integrates geo-contextual information for multimodal generative modeling by constructing datasets and evaluation frameworks.
-3 Method
+Figure 2: The SounDiT framework.
+SounDiT encodes soundscape and scene inputs into a shared space with pre-trained encoders, and performs denoising in latent space with DiT blocks equipped with AdaLN-Zero timestep conditioning.
+Scene information is injected through a Scene Low-Rank Mixer (SLRCM) and a lightweight Scene AdaLN (S-AdaLN) path to enhance geo-contextual consistency, while a Mixture-of-Experts (MoE) Soundscape conditioning module aligns soundscape features with visual landscape tokens.
+3 Methodology
 3.1 Preliminary
 Diffusion Model.
-We adopt a latent-based diffusion model, consisting of a forward diffusion and backward denoising process conducted in the latent space $\mathbf{z}$.
+We adopt a latent-based diffusion model, consisting of a forward diffusion and backward denoising process in the latent space $\mathbf{z}$.
 The forward diffusion $\mathbf{q}$ starts from the initial latent $\mathbf{z}_{0}$, gradually adding Gaussian noise over timestep $T$ based on a fixed variance $\alpha$ and defined as:
-$q(\mathbf{z}_{t}\mid\mathbf{z}_{t-1})=\mathcal{N}\bigl{(}\mathbf{z}_{t};\sqrt{\alpha_{t}}\,\mathbf{z}_{t-1},(1-\alpha_{t})\,\mathbf{I}\bigr{)},\quad t=1,\dots,T.$
+$q(\mathbf{z}_{t}\mid\mathbf{z}_{t-1})=\mathcal{N}\bigl(\mathbf{z}_{t};\sqrt{\alpha_{t}}\,\mathbf{z}_{t-1},(1-\alpha_{t})\,\mathbf{I}\bigr).$ (1)
 The backward denoising process $p_{\theta}$ is trained to predict the less noisy latent:
-$p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t})=\mathcal{N}\bigl{(}\mathbf{z}_{t-1};\,\mu_{\theta}(\mathbf{z}_{t},t,v),\,\Sigma_{\theta}(\mathbf{z}_{t},t,v)\bigr{)},$
+$p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t})=\mathcal{N}\bigl(\mathbf{z}_{t-1};\,\mu_{\theta}(\mathbf{z}_{t},t,v),\,\Sigma_{\theta}(\mathbf{z}_{t},t,v)\bigr).$ (2)
 where $v$ is the conditional signal and $\mu_{\theta}$, $\Sigma_{\theta}$ are predicted through a denoising network $\epsilon_{\theta}(\mathbf{z}_{t},t,v)$.
@@
 To control the influence of the conditioning signal $v$ during backward denoising, we employ classifier-free guidance (CFG) to predict noise $\hat{\boldsymbol{\epsilon}}$ based on both conditional and unconditional signals with specific guidance scale $s$.
-$\hat{\boldsymbol{\epsilon}}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)+s\bigl{[}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,v)-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)\bigr{]}.$
-3.2 Geo-Contextual Soundscape-to-Landscape Generation (GeoS2L) Problem Definition
-In this work, we formulate our task as the Geo-Contextual Soundscape-to-Landscape (GeoS2L) problem, a specialized variant of the audio2image problem, which aims to generate landscape images from environmental soundscapes guided by an optional scene prompt.
-We define a landscape $l$ as the geographical and ecological environment shaped by natural and built features, and a soundscape $s$ as the human-perceived acoustic environment at a given place, comprising both natural and anthropogenic sounds such as bird calls and traffic noise.
-To better support real-world practices, we introduce scene prompt $c$ as an optional semantic label (e.g., park, beach, street) to provide additional geographic semantic context.
-Given a dataset $D=\{\,(s_{i},c_{i},l_{i})\mid i=1,\dots,N\},$ our goal is to learn a generator $G\colon(s_{i},c_{i})\mapsto\hat{l}_{i}$ such that the synthesized image $\hat{l}_{i}$ is both visually realistic and geo-contextually consistent to the ground truth $l_{i}$, quantified by relevance function $R(s,c,l)$.
-This can be formulated as:
-$\mathcal{L}\;=\;\mathbb{E}_{(s_{i},c_{i},l_{i})\sim D}\Bigl{[}\,R(s_{i},c_{i},l_{i})\;-\;R\bigl{(}s_{i},c_{i},\hat{l}_{i}\bigr{)}\Bigr{]}.$
-Minimizing $\mathcal{L}$ enables the model to capture cross-modal correlations among sounding, semantic and visual features, thereby producing landscape images that reflect both the acoustic characteristics of soundscape and intended scene.
-Figure 2: Framework of SounDiT.
-To effectively capture geographic contexts, SounDiT integrates soundscape, scene, and landscape embeddings from pre-trained encoders as inputs.
-The landscape embedding is fused with the timestep embedding and fed into a self-attention layer to extract visual features.
-The scene embedding then guides conditioning via a cross-attention layer.
-The TA-PMoE module further combines these features with soundscape features.
-During inference, a geo-contextual RAG module provides additional geographic contexts to enhance alignment with real-world environmental settings.
+$\hat{\boldsymbol{\epsilon}}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)+s\bigl[\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,v)-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\varnothing)\bigr].$ (3)
+3.2 GeoS2L Task Formulation
+We formulate our task as Geo-Contextual Soundscape-to-Landscape (GeoS2L), extending A2I synthesis with geographic contexts.
+GeoS2L aims to synthesize landscape images from environmental soundscapes guided by an optional scene prompt.
+We define a landscape $l$ as the geographical–ecological environment shaped by natural and built features, a soundscape $s$ as the human-perceived acoustic environment at a place, including natural and anthropogenic sources (e.g., bird calls and traffic noise).
+However, real-world soundscape recordings often lack sufficient information to fully capture the visual characteristics of the surrounding environment.
+For example, a bird call may occur in either rural or urban parks.
+To provide greater flexibility and user control, we incorporate an optional scene prompt $c$ as the semantic geographic context (e.g., park, beach, street).
+This prompt serves as an additional geo-contextual conditioning signal to facilitate generative process toward a specific, context-aware landscape.
+Given a dataset $D=\{(s_{i},c_{i},l_{i})\mid i=1,\dots,N\},$ our goal is to learn a generator $G:\,(s_{i},c_{i})\mapsto\hat{l}_{i}$ such that the synthesized image $\hat{l}_{i}$ is both visually realistic and geo-contextually consistent with the ground truth $l_{i}$, quantified by the relevance function $R(s,c,l)$.
+This can formulated as
+$\mathcal{L}\;=\;\mathbb{E}_{(s_{i},c_{i},l_{i})\sim D}\!\left[\,R(s_{i},c_{i},l_{i})\;-\;R\bigl(s_{i},c_{i},\hat{l}_{i}\bigr)\right].$ (4)
+This enables the model to capture cross-modal correlations among acoustic, semantic, and visual cues, producing landscape images that reflect both the soundscape and the intended scene context.
 3.3 SounDiT
-To support the GeoS2L task, we propose SounDiT, a diffusion-based generative model that synthesizes geographically coherent landscape images conditioned on environmental soundscapes and optional scene prompts (Figure 2).
-SounDiT advances geo-contextual landscape generation through two key innovations: (1) the SounDiT Block, which integrates sound, scene, and visual cues into the denoising process; and (2) a geo-contextual Retrieval-Augmented Generation (RAG) to guide image generation with additional geographic contexts at inference time.
+To effectively and efficiently inject soundscape and scene context, each SounDiT block introduces (1) a Mixture-of-Experts (MoE) Soundscape Conditioning module, (2) a Scene Low-Rank Mixer (SLRCM) module, and a Scene AdaLN (S-AdaLN) module to improve scene consistency.
 Multimodal Encoders.
-To incorporate acoustic, semantic and visual features into SounDiT, we employ three encoder modules: (1) a Soundscape Encoder $\mathcal{E}_{\mathrm{Sound}}$, (2) a Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ and (3) a Landscape Encoder–Decoder pair $\bigl{(}\mathcal{E}_{L},\mathcal{E}^{\prime}_{L}\bigr{)}$.
-The Soundscape Encoder $\mathcal{E}_{Sound}$ embeds soundscape $s$ into an audio embedding $e_{\mathrm{sound}}$.
-The Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ encodes the optional scene prompt $c$ into a text embedding $e_{\mathrm{scene}}$.
-The Landscape Encoder $\mathcal{E}_{L}$ compresses a landscape image $l_{i}\in\mathbb{R}^{H\times W\times 3}$ into a latent embedding $e_{\mathrm{landscape}}$, and the Landscape Decoder $\mathcal{E^{\prime}}_{L}$ reconstructs the latent embedding back into a landscape image $\hat{l_{i}}\in\mathbb{R}^{H\times W\times 3}$.
+Following latent diffusion models (LDM) [66], landscape images are reconstructed within the latent space.
+Given a landscape image $l\!\in\!\mathbb{R}^{H\times W\times 3}$, a VAE encoder–decoder $(\mathcal{E}_{L},\mathcal{D}_{L})$ is employed to encode it into the latent embedding $\mathbf{e}_{l}=\mathcal{E}_{L}(l)\in\mathbb{R}^{h\times w\times c}$ and reconstruct the latent embedding back into a image $\hat{l}=\mathcal{D}_{L}(\mathbf{e}_{l}).$ To integrate acoustic features and geo-contextual semantics into denoising process, a multi-modal encoder is adopted to encode the soundscapes $s$ and the scene prompt $c$ into a shared latent space, resulting in a soundscape embedding $\mathbf{e}_{s}$ and a scene embedding $\mathbf{e}_{c}$.
 SounDiT block.
-Despite the demonstrated effectiveness of DiT in image generation, they have not been integrated into audio2image tasks.
-To effectively model geographic contexts from both soundscape and scene inputs, we propose the SounDiT block, a geo-contextual DiT to generate landscape images that align with intended geographic settings.
-The SounDiT operates in three stages: First, the timestep embedding $e_{\mathrm{t}}$ is fused with landscape latent embedding $e_{\mathrm{landscape}}$, and fed into a multi-head self-attention layer to capture visual features.
-Second, a multi-head cross-attention conditioned on the scene embedding $e_{\mathrm{scene}}$ is applied to enrich these features with scene semantics, producing scene-landscape representation.
-Third, to further integrate multi-level soundscape features, we introduce Time-Aware Prototype Mixture-of-Experts (TA-PMoE) module, composed of $M$ parallel cross-attention layers as experts.
-Each cross-attention layer fuses the scene-landscape representation with the soundscape embedding $e_{\mathrm{sound}}$, resulting in $M$ soundscape-scene-landscape representations $\mathbf{P}=[\,\mathbf{p}_{1},\dots,\mathbf{p}_{M}\,]$.
-A router $\mathbf{w}$ then calculates a weighted combination of these representations $\mathbf{P}$ based on temperature-scaled Softmax:
+The SounDiT block operates in four stages.
+First, we apply multi-head self-attention layer conditioned by the timestep embedding $e_{\mathrm{t}}$ via AdaLN-Zero [26], preserving compatibility with pretrained DiT backbones.
+Second, we design a lightweight Scene Low Rank Context Mixer (SLRCM) module to inject scene context.
+SLRCM introduces a low-rank residual path inside each SounDiT block.
+Given an input token $x\in\mathbb{R}^{B\times N\times D}$ and scene embedding $\mathbf{e}_{c}\in\mathbb{R}^{B\times D_{c}}$, SLRCM modulates tokens through a rank-$r$ linear operator parameterized by $\mathbf{e}_{c}$.
+We specify the low-rank content mixer $A(\mathbf{e}_{c})$ that transforms token features along an $r$-dimensional path, together with a sample-wise strength $s(\mathbf{e}_{c})$ that controls the overall contribution of this path:
+$A(\mathbf{e}_{c})=W_{q}\,\mathrm{Diag}\!\big(\tanh(\phi(\mathbf{e}_{c}))\big)\,W_{v},$ (5)
+where $W_{q}\!\in\!\mathbb{R}^{D\times r}$ and $W_{v}\!\in\!\mathbb{R}^{r\times D}$ are learned low-rank projections.
+The mapping $\phi:\mathbb{R}^{D_{c}}\!\to\!\mathbb{R}^{r}$ converts the scene embedding into an $r$-dimensional gating vector, and the diagonal operator applies this gate element-wise along the rank-$r$ channel, yielding a compact, diagonally gated low-rank transform.
+In addition, we define a positive, sample-wise strength to regulate the residual magnitude:
+$s(\mathbf{e}_{c})=g(\mathbf{e}_{c})\,\mu\,\tanh(\alpha),$ (6)
+where $g:\mathbb{R}^{D_{c}}\!\to\!\mathbb{R}_{>0}$ produces a per-sample scale (e.g., via softplus), $\mu$ is a global guidance scalar, and $\alpha$ is a learnable scalar bounded.
+Given these definitions, the token update becomes:
+$x^{\prime}\;=\;x\;+\;s(\mathbf{e}_{c})\,\mathrm{LN}(x)\,A(\mathbf{e}_{c}),$ (7)
+where $\mathrm{LN}(\cdot)$ denotes layer normalization.
+SLRCM thus provides a compact, diagonally gated low-rank modulation that preserves the pre-trained attention structure while injecting scene priors with low computational costs.
+Third, to further integrate multi-level soundscape features, we introduce MoE Soundscape Conditioning module, composed of $M$ experts that share the key $K_{s}$ and value $V_{s}$ while maintaining expert-specific, low-rank query $Q$.
+Given input token $x\in\mathbb{R}^{B\times N\times D}$,we compute $K_{s}=s\,W_{K},V_{s}=s\,W_{V}$ once and for each expert $m\in\{1,\dots,M\}$ form a low-rank query $Q_{m}$
+$Q_{m}=f(x)\,W^{(m)}_{q\downarrow}W^{(m)}_{q\uparrow}\!$, (8)
+where $f$ is a per-token Layer Norm. Each expert then produces
+$Z_{m}\;=\;\mathrm{MHA}\bigl(Q_{m},\,K_{s},\,V_{s}\bigr)\,W_{O},$ (9)
+where $W_{O}$ is shared across experts, $Z_{m}$ refers to soundscape-conditioned token features, and $MHA$ is multi-head attention.
+Routing weights are computed by a temperature-scaled dot product between the audio summary and learnable prototypes $\mathbf{P}\in\mathbb{R}^{d\times M}$ with time augmentation introduced through a projected timestep embedding:
+$w=\operatorname{Softmax}\!\Bigl(\tfrac{1}{\tau}\;\mathbf{e}_{s}^{\top}\bigl(\mathbf{P}+\mathbf{W}_{t}\,e_{t}\,\mathbf{1}^{\top}\bigr)\Bigr),$ (10)
+where $e_{t}$ is the timestep embedding projected by $\mathbf{W}_{t}$.
+Finally, we aggregate the experts via a top-$k$ soft mixture and apply a global audio gate to update tokens in each SounDiT block:
+$x^{\prime}\;=\;x\;+\;\gamma\sum_{m\in\mathcal{K}}\operatorname{softmax}\!\bigl(w_{m}/\tau_{m}\bigr)_{m}\;Z_{m},$ (11)
+where $\gamma=\tanh(\mathbf{e}_{s})$ is a scalar gate, yielding a smooth, zero-centered bounded activation.
+By enabling expert specialization under a fixed K/V budget, the MoE provides scalable conditioning that improves geo-contextual consistency across diverse soundscapes.
+Finally, to enhance geo-contextual consistency, we introduce a practical-informed S-AdaLN, an AdaLN-Zero extension, which derives scale–shift parameters from the timestep embedding $\mathbf{e}_{t}$ and the scene embedding $\mathbf{e}_{c}$ via learnable bounded mixing.
+The modulated tokens are processed by a pointwise feed-forward network, followed by a gated residual add, yielding the block output for noise-residual prediction.
+By hierarchically fusing visual, scene, and soundscape cues, SounDiT generates geo-contextually coherent landscapes.
+4 Datasets
+Table 1: Comparison of audio-visual datasets. $\dagger$} denotes partial/limited geo-context.
+Dataset Modalities # Size Geo-Contextual Viewpoint
+ACAV100M [44] Video 100M - Ground
+VEGAS [87] Video 28K - Ground
+VGGSound [10] Video 200K $\dagger$} Ground
+IntoTheWild [46] Video 142 Ground
+Landscape [45] Video 1747 Ground
+SoundingEarth [30] Audio & Image 50K Overhead
 […]
-Here, $\tau>0$ controls the sharpness of the distribution: smaller values of $\tau$ concentrate weight on the most relevant soundscape-scene-landscape representation, while larger values lead to a smoother allocation across representations.
-Finally, the combined soundscape-scene-landscape representation is passed through a pointwise feedforward network to produce the denoised output.
-By hierarchically fusing visual, scene, and soundscape representations, SoundDiT effectively generates geo-contextually coherent landscape images.
-Geo-contextual RAG Inference.
-In practice, additional geo-contextual information, such as location, scene and spatial layout [47], may be collected alongside soundscapes and landscapes to guide image generation.
-Thus, we propose a flexible geo-contextual RAG pipeline to further enhance the geographic coherence of the generated landscape images.
-Our method consists of two stages: (1) Geo-contextual Knowledge Database (GeoKD), where we store soundscape embeddings along with associated scene categories and layout representations with geographic locations; and (2) Geo-contextual Initialization, where the most relevant scene and layout are retrieved based on soundscape similarity and geographic proximity.
-These retrieved cues are then used to initialize the diffusion generation to enhance geographic coherence.
-4 Datasets: SoundingSVI and SonicUrban
-Prior audio2image studies have primarily relied on general datasets such as VEGAS [85], VGGSound [9], and ACAV100M [40].
-Despite their proven effectiveness, these datasets lack geographic contexts.
+[…]
+Prior A2I studies have primarily relied on large-scale, general-domain audio-visual datasets, such as VEGAS [87], VGGSound [10], and ACAV100M [44].
+These datasets primarily focus on sound-source localization and general audio-visual correspondence learning, lacking the specific geographical context required for soundscape-to-landscape generation.
 For instance, geographers may be more interested in generating a wetland implied by a frog call, rather than an image of the frog itself.
@@
 To address this gap, we introduce two new multimodal datasets: SoundingSVI and SonicUrban, designed to support GeoS2L generation with large-scale, geographically diverse, and context-rich data.
-Figure 3 provides several sample images of these two datasets.
-[…]
-Dataset Modalities # Size Geo-Contextual Scene Diversity Spatial Coverage Viewpoint
-ACAV100M [40] Video 100M -- -- -- Ground-level
-VEGAS [85] Video 28K -- -- -- Ground-level
-[…]
-IntoTheWild [42] Video 142 $\checkmark$ Low -- Ground-level
-Landscape [41] Video 1747 $\checkmark$ Medium -- Ground-level
-SoundingEarth [28] Audio & Image 50K $\checkmark$ High High Overhead
-SoundingSVI (Ours) Audio & Image 330K $\checkmark$ High High Ground-level
-SonicUrban (Ours) Video 250K $\checkmark$ High High Ground-level
-Figure 3: Representative samples from the SoundingSVI and SonicUrban datasets, showing diverse geographic contexts.
 SoundingSVI.
@@
 A sound source localization model [61] was further applied to match each clip with the most relevant landscape image.
-In total, SoundingSVI dataset comprises 334,495 soundscape-landscape pairs in 90 countries.
-To support optional scene prompts during training, a scene type for each soundscape-landscape pair was predicted by a scene classification model [27].
+To ensure temporal consistency and account for environmental changes, we refined the dataset by discarding soundscape-SVI pairs with significant time gaps.
+Furthermore, to obtain annotated place semantics, a Vision-Language Model (VLM), Qwen2.5-VL-7b, [2] was used to annotate each pair with a scene context (e.g., street and residential neighborhood), which serve as the scene prompts for our model.
+In total, SoundingSVI dataset comprises 169,221 soundscape-landscape pairs in 90 countries.
 SonicUrban.
@@
 The same human voice detection model [74] and sound source localization model [61] used in SoundingSVI were applied to filter human-voice dominated clips and identify the most representative frame for each audio clip, respectively.
-In total, the constructed SonicUrban consists of 250,310 soundscape-landscape pairs across 131 cities and 97 countries.
-A scene classification model [27] was further employed to generate a scene prompt for each pair as a geographic scene condition.
-5 Geo-Contextual Evaluation Framework
-Figure 4: Geo-contextual evaluation framework.
-The ground truth and output landscape images are assessed at the element-, scene-, and human-perception levels following geography and urban planning practices.
-Prior audio2image studies commonly rely on evaluation metrics such as Fréchet Inception Distance (FID) [30], Audio-Image Similarity (AIS) [30], and Image-Image Similarity (IIS) [5], to assess visual or audio fidelity, overlooking the underlying geographic settings and target scenes.
+In total, the constructed SonicUrban consists of 236,674 soundscape-landscape pairs across 131 cities and 97 countries.
+The same VLM [2] was further employed to generate a scene prompt for each pair as a geographic scene condition.
+5 Geo-Contextual Evaluation Suite
+Prior A2I studies commonly rely on evaluation metrics such as Fréchet Inception Distance (FID) [33], Audio-Image Similarity (AIS) [33], and Image-Image Similarity (IIS) [5], to assess visual or audio fidelity.
+However, such metrics do not assess whether the generated image is coherent with the underlying geographic setting or the intended target scene.
 Since soundscapes and landscapes co-occur in the same space, they are expected to share similar environmental characteristics and place settings.
-Following this hypothesis, we introduce a novel practically-informed geo-contextual evaluation framework — Place Similarity Score (PSS) — that measures the environmental settings of soundscape and landscape across element, scene, and human perception levels (Figure 4).
+Following this hypothesis, we introduce Place Similarity Score (PSS), a practically informed geo-contextual evaluation that quantifies the underlying place setting reflected in images across the element-, scene-, and human-perception levels.
 These three levels of measurements of environmental characteristics have been widely employed in geographic and urban planning practices [53, 41].
-Table 2: Place Similarity Score (PSS)
-Score Definition
-$\text{PSS}_{\text{element}}$ $\frac{1}{n}\sum_{i=1}^{n}\frac{E_{i}\cdot\hat{E}_{i}}{\|E_{i}\|\|\hat{E}_{i}\|}$
-$\text{PSS}_{\text{scene}}$ $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{P_{i}^{\text{k}}\cap T_{i}^{\text{k}}\neq\varnothing\right\}$
-$\text{PSS}_{\text{perception}}$ $\frac{1}{n}\sum_{i=1}^{n}\left\|R(l_{i})-R(\hat{l}_{i})\right\|_{1}$
-Element-level Evaluation quantifies the similarity of geographic elements (e.g., trees [43], sky [23], water bodies [69], traffic signs [3], and buildings [13]) between the geographical contexts of the input soundscape and generated landscapes [35].
-We employ the DeepLabV3 semantic segmentation model [12], pre-trained on the ADE20K dataset [84], to identify 150 predefined geographic elements $E_{i}$ in each image $i$.
-For each image $i$, we calculate a normalized element distribution vector $E_{i}=\{e_{i}^{1},e_{i}^{2},\dots,e_{i}^{k}\}$, where each $e_{i}^{k}$ represents the ratio of pixels identified as the $k$-th element.
-[17].
-We calculate element ratio vectors $E_{i}$ and $\hat{E_{i}}$ from the ground truth landscape image $l_{i}$ and their corresponding image $\hat{l_{i}}$.
-We then measure the cosine similarity between each pair of real and generated images and report the average across all samples $n$ as the $PSS_{element}$.
-Scene-level Evaluation is designed to evaluate the consistency of the entire environmental scene (e.g., “forest,” “beach,” or “residential area”) between the generated landscape images and their corresponding ground truth images [35].
+Element-level PSS.
+This metric measures the agreement in geographic elements (e.g., trees [47], sky [24], water bodies [70], traffic signs [3], and buildings [14]) between the ground-truth and generated landscapes [40].
+We use DeepLabV3 [13] pre-trained on ADE20K [86] to segment K=150K\!=\!150 predefined elements.
+For each image $i$, let $\mathbf{e}_{i}\in\mathbb{R}^{K}$ and $\hat{\mathbf{e}}_{i}\in\mathbb{R}^{K}$ be the normalized element-ratio vectors derived from the ground-truth image $l_{i}$ and the generated image $\hat{l}_{i}$, respectively.
+Formally, The element-level score is calculated as:
+$\mathrm{PSS}_{\text{elem}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\mathbf{e}_{i}^{\top}\,\hat{\mathbf{e}}_{i}}{\|\mathbf{e}_{i}\|_{2}\,\|\hat{\mathbf{e}}_{i}\|_{2}},$ (12)
+where $\mathbf{e}_{i}=(e_{i1},\dots,e_{iK})^{\top}$ and $\hat{\mathbf{e}}_{i}=(\hat{e}_{i1},\dots,\hat{e}_{iK})^{\top}$; $n$ is the number of evaluation images.
+Scene-level PSS.
+This metric is designed to evaluate the consistency of the entire environmental scene (e.g., forest, beach, or residential area) between the generated landscape images and their corresponding ground truth images [40].
 To infer scene categories, we employ a ResNet50 model [29], pre-trained on the Places365 dataset [85], to identify 365 predefined scene categories.
 For each soundscape $s$, we predict the top-$k$ scene labels and evaluate whether the generated image $\hat{l_{i}}$ aligns with the same scene category as its ground truth image $l_{i}$.
-The $PSS_{scene}$ is then calculated by its top-$k$ accuracy, where $k=1$ or $5$.
+$\mathrm{PPS}_{\text{Scene}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{P_{i}^{\text{k}}\cap T_{i}^{\text{k}}\neq\varnothing\right\},$ (13)
+where $k=1$ or $5$.
 $n$ is the total number of images, $P_{i}^{\text{k}}$ denotes the top-$k$ predicted category set for the $i$-th generated image, $T_{i}^{\text{k}}$ denotes the top-$k$ ground truth category set for the $i$-th real image.
-Human Perception-level Evaluation is designed to evaluate whether the generated and real landscape images evoke similar human subjective perceptions and feelings, such as whether a place makes people feel “safe” or “lively” [82].
-We utilize a DenseNet121 model [31], pre-trained on the MIT Place Pulse dataset [55], to measure six dimensions of human perceptions of environment, including “Safe,” “Beautiful,” “Depressing,” “Lively,” “Wealthy,” and “Boring”.
-For each image $i$, we compute the perception scores in six dimensions for the generated image $\mathrm{R}(\hat{l}_{i})$ and its corresponding ground truth $\mathrm{R}(l_{i})$.
-$PSS_{perception}$ is then defined as the average L1 distance across six dimensions, where $n$ is the number of the test set, $R$ is the Place Perception scores of each dimension.
-By integrating metrics across element-, scene-, and human perception levels (Table 2), we provide a holistic geo-contextual evaluation framework.
+Human Perception-level PSS.This metric is designed to evaluate whether the generated and real landscape images evoke similar human subjective perceptions and feelings, such as whether a place makes people feel “safe” or “lively” [83].
+We utilize a DenseNet121 model [34], pre-trained on the MIT Place Pulse dataset [58], to measure six dimensions of human perceptions of environment, including safe, beautiful, depressing, lively, wealthy, and boring.
+For each image $i$, we compute six-dimensional place-perception scores for the ground-truth image $l_{i}$ and the generated image $\hat{l}_{i}$.
+The perception-level PSS is defined as:
+$\mathrm{PSS}_{\text{perc}}=\frac{1}{n}\sum_{i=1}^{n}\left\|\mathbf{R}(l_{i})-\mathbf{R}(\hat{l}_{i})\right\|_{1},$ (14)
+where $\mathbf{R}(\cdot)\in\mathbb{R}^{6}$ denotes the six-dimensional perception score vector, and $n$ is the number of test images.
+By integrating metrics across element-, scene-, and human perception levels, we provide a holistic geo-contextual evaluation framework.
 Informed by geographic domain knowledge and urban planning practices, we extend beyond measuring visual quality to assessing whether generated landscape images are contextually aligned with the environmental characteristics in input soundscapes.
-6 Experiments
-Figure 5: Visual comparison of landscape images generated by baseline models (CoDi, Sound2Scene, AudioToken Pre-trained/SD1/SD2, GlueGen) and our proposed SounDiT across SoundingSVI, SonicUrban, and two benchmark datasets.
-Ground truth landscape images are provided in the last column.
-Our proposed SounDiT captures significantly more geographic contexts than baselines.
+Table 2: Quantitative evaluation results across SounDiT and baseline models on two proposed datasets and benchmark datasets.
+$\dagger$} indicates pre-trained models.
+We compare our SounDiT model with baselines using general metrics (FID, AIS, IIS) and our PSS metrics.
+General Metrics Place Similarity Score (PSS)
+[…]
+[…]
+Sound2Scene [72] 78.137 0.586 0.623 0.467 0.374 0.735
+[…]
+AudioToken (SD1) [81] 105.472 0.509 0.572 0.527 0.553 0.797
+AudioToken (SD2) [81] 149.473 0.504 0.559 0.249 0.311 0.845
+GlueGen [63] 74.250 0.500 0.566 0.400 0.461 0.776
+PixArt + MHCA [12] 34.108 0.518 0.578 0.474 0.390 0.743
+[…]
+[…]
+Sound2Scene [72] 49.681 0.503 0.498 0.332 0.324 0.829
+[…]
+AudioToken (SD1) [81] 95.210 0.505 0.503 0.419 0.484 0.979
+AudioToken (SD2) [81] 165.014 0.502 0.504 0.298 0.455 0.980
+GlueGen [63] 64.434 0.498 0.499 0.405 0.481 0.805
+PixArt + MHCA [12] 41.456 0.517 0.592 0.411 0.396 0.796
+[…]
+6 Experiment
 6.1 Implementation Details
@@
 We adopt the pre-trained Variational Autoencoder (VAE) modules from Latent Stable Diffusion [66] as our landscape encoder–decoder pair, which were originally trained on the COCO dataset [49].
-For soundscape conditioning, we employ the ImageBind-Huge audio encoder [22], pre-trained on two million AudioSet clips [21] covering a wide range of environmental soundscapes.
-We integrate a scene encoder based on the T5 architecture [62], which transforms high-level semantic scene prompts into latent embeddings to provide geographic semantics.
-Our SounDiT Block is initialized with the weights of PixArt-$\alpha$, a pre-trained text-to-image DiT model [10], and four experts are employed in TA-MoE Module.
-All components are trained using a learning rate of $1\times 10^{-4}$.
-Experiments are conducted on NVIDIA A100 and A6000 GPUs.
-Benchmark Datasets.
-We curated two additional GeoS2L benchmark datasets to evaluate model generalizability.
-The first is a geo-contextual subset of VGGSound [9], comprising 16 classes with environmental semantics.
-It contains 6,713 soundscape–landscape pairs.
-The second combines the Landscape [41] and IntoTheWild [42] datasets, resulting in 1,889 soundscape-landscape pairs.
+For soundscape and scene prompt conditioning, we employ the ImageBind-Huge [23], pre-trained on two million AudioSet clips [21] covering a wide range of environmental, chosen for its accuracy–efficiency balance among several encoders[78].
+Our SounDiT is trained using a learning rate of $1\times 10^{-4}$.
+Experiments are conducted on NVIDIA H100, A100 and A6000 GPUs.
+During training, we set the soundscape guidance scalar to $\mu=1.0$.
+For the scene condition, we employ a learnable scaling parameter $\alpha$ initialized to 0 to stabilize early training by starting from an identity mapping.
+At inference stage, we apply classifier-free guidance with a scale of $4.0$ for both the soundscape and the scene prompt.
 Baseline Models.
-We selected four state-of-the-art audio2image models, including CoDi [72], Sound2Scene [70], AudioToken [80], and GlueGen [61] as baseline models.
+We compare against state of the art audio to image models: CoDi [73], Sound2Scene [72], AudioToken [81], GlueGen [63] and a PixArt [12] variant augmented with multihead cross attention for the soundscape condition (PixArt+MHCA).
 For CoDi, we utilized its official pre-trained model.
 For AudioToken, we employed the pre-trained model without additional training, and trained it on our datasets using Stable Diffusion 1 and 2 as its backbone, respectively.
-The other two baselines are training from scratch based on their released code and training parameters.
+The other baselines are training from scratch based on their released code and training settings.
 Benchmark Evaluation Metrics.
-To evaluate the model performance, we adopted FID [30], AIS [30], and IIS [5], effective in previous audio-conditioned image generation studies [5, 71].
+To evaluate the model performance, we adopted FID [33], AIS [33], and IIS [5], effective in previous audio-to-image generation studies [5, 71].
 FID compares Gaussian distributions of features extracted from real and generated images, using a pre-trained Inception model.
@@
 6.2 Main Results
+Figure 3: Visual comparison of landscape images generated by baseline models (CoDi, Sound2Scene, AudioToken(AT), GlueGen, PixArt) and our proposed SounDiT across SoundingSVI and SonicUrban.
+Ground truth landscape images are provided in the last column.
+Our proposed SounDiT significant improvements over baseline models in generating geographical context.
 […]
-General Metrics Place Similarity Score
+Table 3: Component ablations on SoundingSVI (169K) with two experts in the MoE Soundscape Conditioning module.
+We assess the individual contributions of Scene Low-Rank Content Mixer (SLRCM) and the S-AdaLN.
 […]
-[…]
-Sound2Scene 78.137 0.586 0.623 0.467 0.374 0.735
-[…]
-AudioToken (SD1) 105.472 0.509 0.572 0.527 0.553 0.797
-AudioToken (SD2) 149.473 0.504 0.559 0.249 0.311 0.845
-GlueGen 74.250 0.500 0.566 0.400 0.461 0.776
-SounDiT (Ours) 28.173 0.542 0.663 0.652 0.578 0.728
-[…]
-Sound2Scene 49.681 0.503 0.498 0.332 0.324 0.829
-[…]
-AudioToken (SD1) 95.210 0.505 0.503 0.419 0.484 0.979
-AudioToken (SD2) 165.014 0.502 0.504 0.298 0.455 0.980
-GlueGen 64.434 0.498 0.499 0.405 0.481 0.805
-SounDiT (Ours) 33.246 0.517 0.592 0.465 0.495 0.788
-[…]
-Sound2Scene 140.244 0.617 0.741 0.479 0.640 0.769
-[…]
-AudioToken (SD1) 138.785 0.585 0.695 0.364 0.534 0.846
-AudioToken (SD2) 118.536 0.637 0.740 0.512 0.724 0.855
-GlueGen 195.625 0.584 0.645 0.362 0.582 0.815
-SounDiT (Ours) 155.471 0.628 0.744 0.521 0.741 0.757
-[…]
-Sound2Scene 122.460 0.600 0.596 0.287 0.378 0.770
-[…]
-AudioToken (SD1) 163.540 0.532 0.555 0.207 0.311 0.789
-AudioToken (SD2) 157.706 0.583 0.587 0.251 0.404 0.803
-GlueGen 156.777 0.536 0.547 0.233 0.249 0.839
-SounDiT (Ours) 113.111 0.575 0.588 0.336 0.392 0.767
-Figure 5 presents landscape images generated by our proposed SounDiT and baseline models across SoundingSVI, Sonic Urban, and two benchmark datasets.
-The results generated by SounDiT show greater geo-contextual similarity to ground truth images, highlighting potential applications of SounDiT in geography and urban planning.
-To evaluate model performance, we adopt general metrics (FID, AIS, IIS) in addition to our proposed geo-contextual PSS framework.
-As reported in Table 3, SounDiT consistently outperforms baselines on both SoundingSVI and SonicUrban datasets, with notable FID improvements (74$\rightarrow$28 and 49$\rightarrow$33), and achieves superior or competitive results across all metrics.
-SounDiT consistently ranks highest under PSS, indicating its ability to ensure geographic coherence in generated images.
-These results validate the effectiveness of PSS in measuring geo-contextual coherence beyond general metrics.
-The performance drop of SounDiT on the smaller IntoTheWild&Landscape dataset further emphasizes the crucial needs for large-scale training datasets such as SoundingSVI and SonicUrban.
+Full Model 19.195 0.538 0.750 0.734
+w/o SLRCM + S-AdaLN 25.375 0.511 0.539 0.428
+w/o SLRCM 20.335 0.534 0.728 0.704
+w/o S-AdaLN 23.435 0.529 0.629 0.572
+Practical Applications.
+To enhance real-world utility and user control, SounDiT supports scene-conditioned generation.
+Given a fixed soundscape, modifying the scene prompt generates geo-contextual, semantically coherent images that preserve both auditory information and scene context (Fig.
+4).
+This supports practical applications such as soundscape-guided urban design, informing evidence-based design strategies to promote public health, safety, and environmental comfort.
+First, SounDiT can support soundscape-guided urban design by generating plausible visual place contexts from a target soundscape (quiet/restorative or a recorded clip), helping designers plan actionable elements such as greenery and water.
+Second, SounDiT facilities noise mitigation strategies by visualizing the built environment, enabling diagnosis of noise sources beyond decibel (dB), informing interventions (green buffers, traffic calming) for planning decisions.
+Moreover, SounDiT enables inclusive, accessible design by translating soundscape into visual landscapes, broadens participation in urban design/planning practices, especially for hearing-impaired participants.
+Figure 4: Scene-conditioned generation. With the same soundscape input, SounDiT generates visually distinct yet acoustically consistent images.
 User Study.
 In addition to these quantitative metrics, we further conducted a user study to assess whether the generated landscape images align with human perceptions of geographic contexts.
-Specifically, participants were asked to complete two matching tasks: (1) selecting the generated image that best aligned with the given soundscape; and (2) identifying which generated landscape image most closely matched a ground-truth landscape.
+Specifically, 17 participants were asked to complete two matching tasks: (1) selecting the generated image that best aligned with the given soundscape; and (2) identifying which generated landscape image most closely matched a ground-truth landscape.
 The average matching accuracy is 86.13%, indicating a strong perceptual alignment between soundscapes and their corresponding generated images.
 This demonstrates the effectiveness of SounDiT for GeoS2L.
+Table 4: Expert scalability of MoE cross-attention on SoundingSVI (169K). We vary the number of audio experts $M\!\in\!\{2,4,6,8\}$ while keeping other settings fixed.
+Experts 2 4 6 8
+[…]
+$\mathrm{PSS}_{\mathrm{Scene}}\uparrow$ 0.734 0.741 0.742 0.753
 6.3 Ablations
-Table 4: Ablation results. We assess the contribution of key components to the overall performance of SounDiT.
-[…]
-TA-MoE 28.173 0.652
-MHCA 34.108 0.602
-(a) SounDiT Block Design
-[…]
-w/ RAG 0.652 0.578
-w/o RAG 0.479 0.514
-(b) Geo-contextual RAG
-We conducted the ablation study on SoundingSVI.
-First, we replace the TA-PMoE module with a standard multi-head cross-attention layer, forming SounDiT-MHCA.
-Second, we disable the RAG module in SounDiT-TA-MoE, yielding SounDiT-TA-MoE w/o RAG.
-As reported in Table 4(a), SounDiT-TA-MoE achieves a lower FID and a higher $PSS_{element}$, indicating that the expert routing (TA-MoE) is essential for modeling multi-level soundscape and visual features, thereby preserving visual quality and geographic coherence in the generated images.
-Table 4(b) further shows that incorporating geo-contextual RAG inference improves both element- and scene-level PSS of generated landscape images, highlighting the benefits of integrating geographic contexts during inference.
-7 Concluding Remarks
+To validate our architectural design and quantify the impact of our proposed modules of SounDiT, we conduct two types of ablation experiments: component ablations and component scalability ablations.
+First, to assess the contribution of scene prompt and two scene conditioning modules, we perform experiments by first removing both SLRCM and S-AdaLN, and then removing each module individually.
+As shown in Table 3, removing either module results in a noticeable drop in performance.
+These results indicate that injecting scene condition both before and after the MoE Soundscape Conditioning module improves geo-contextual alignment and overall quality.
+Second, we test the scalability of MoE Soundscape Conditioning by varying the number of experts within this module (Table 4).
+Results show that the model with more experts achieves better performance in FID and scene consistency.
+7 Conclusion
 In this work, we present Geo-Contextual Soundscape-to-Landscape (GeoS2L), a practically-informed novel generation problem that synthesizes realistic landscape images from environmental soundscapes.
-We explicitly integrate geographic domain knowledge into multimodal generative modeling.
 To support GeoS2L, we construct SoundingSVI and SonicUrban, two large-scale, multi-modal, geo-contextual datasets, capturing diverse sounding and visual environments.
-We propose SounDiT, a novel diffusion-based framework that captures geographic semantics embedded in sounding environments.
+We propose SounDiT, a scalable diffusion-based framework that captures both sounding environment and scene context by MoE Soundscape Conditioning and Scene Low-Rank Content Mixer.
 We further design a novel geo-contextual evaluation framework, the Place Similarity Score (PSS), which assesses geographic and environmental consistency.
@@
 The proposed geo-contextual datasets, model, and evaluation framework effectively integrate geographic knowledge into generative modeling, establishing a solid benchmark for the GeoS2L task.
-Limitations.
-While our work introduces a novel problem, GeoS2L, several limitations remain.
-First, the SoundingSVI dataset relies on geotagged audio from Aporee, which is a crowd-sourced platform that may contain spatial inaccuracies.
-Future efforts may focus on refining spatial alignment and improving dataset quality through more robust data processing procedures.
-Second, the current PSS evaluation metrics only consider three practically-informed levels.
-Future studies could integrate more geographic domain-specific knowledge to support geo-contextually coherent image generation.
+Acknowledgements
+We acknowledge ACCESS, supported by the U.S.
+National Science Foundation, for allocation SOC250002.
+We also acknowledge resources from the NAIRR Pilot (NAIRR250075).
+We acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing HPC resources.